OpenStudy (anonymous):

Sin^-1(sin 4pi/7) Ive been on this for the past 30 mins. I just cant understand it

OpenStudy (***[isuru]***):

here is an example for u \[\sin \frac{ \pi }{ 6 } = \frac{ 1 }{ 2 }\] then \[\sin^{-1} \frac{ 1 }{ 2 } = \frac{ \pi }{ 6 }\] now multiply both sides by (sin) \[\sin [\sin^{-1} \frac{ 1 }{ 2 } ] = \sin \frac{ \pi }{ 6 }\] but \[\sin \frac{ \pi }{ 6 } = \frac{ 1 }{ 2 }\] there for \[\sin [\sin^{-1} \frac{ 1 }{ 2 } ] = \ \frac{ \ 1 }{ 2 }\] did u get it ?

OpenStudy (anonymous):

I understand the concept, but i dont understand what sin\[\frac{ 4\pi }{ 7 }\] is equal to

OpenStudy (anonymous):

How do you find that out?

OpenStudy (***[isuru]***):

u want to find sin (4pi/7) ?

OpenStudy (anonymous):

Yup.